OpenStudy (anonymous):

if f(x) = x/1+x and g(x)= sin 3x what is f(g(x)) and what is its domain

OpenStudy (anonymous):

Do you understand how to find the composite? Basically, you are exchanging g(x) for every X in the F(x) domain is where the denominator is equal to zero Post the composite and we can go from there

OpenStudy (anonymous):

that is what I thought so I had f(g(x)) = sin3x/1+sin3x but when I plugged that into my homework it said that it was incorrect

zepdrix (zepdrix):

Was the problem presented like this: \(\Large\rm f(x)=\dfrac{x}{1+x}\) or like this: \(\Large\rm f(x)=\dfrac{x}{1}+x\) I'm assuming it was the first one. Your solution looks correct, you probably just need to format it properly. f(g(x)) = sin(3x)/(1+sin(3x))

OpenStudy (anonymous):

Your answer is correct, but I think they are requiring parenthesis...you should put what is in the denominator in parenthesis...it can be read a couple of ways...see example above

OpenStudy (anonymous):

okay thank y'all very much!

OpenStudy (anonymous):

@zepdrix the first way for f(x) is correct but then how would I go about figuring out f(f(x))

zepdrix (zepdrix):

\[\Large\rm f(\color{royalblue}{x})=\frac{\color{royalblue}{x}}{1+\color{royalblue}{x}}\]So we're going to take our entire f(x) and plug it in for our x's.\[\Large\rm f(\color{royalblue}{f(x)})=\frac{\color{royalblue}{f(x)}}{1+\color{royalblue}{f(x)}}\]

zepdrix (zepdrix):

And so on the right side, we'll plug in the equation for our f(x)'s, yes?

OpenStudy (anonymous):

yes so you'll have x/1+x all over 1+ (x/1+x)

zepdrix (zepdrix):

Mmm yes, good good :) \[\Large\rm f(\color{royalblue}{f(x)})=\frac{\color{royalblue}{\left(\frac{x}{1+x}\right)}}{1+\color{royalblue}{\left(\frac{x}{1+x}\right)}}\]

zepdrix (zepdrix):

Again though if you want to put that into the website, you'll need to use A LOT more brackets lol

zepdrix (zepdrix):

Might be worth your time to simplify first though.

OpenStudy (anonymous):

and to simplify I would multiply the top and bottom by (1-x)?

zepdrix (zepdrix):

It will work out better to multiply top and bottom by (1+x)

zepdrix (zepdrix):

\[\Large\rm f(\color{royalblue}{f(x)})=\frac{\color{royalblue}{\left(\frac{x}{1+x}\right)}}{1+\color{royalblue}{\left(\frac{x}{1+x}\right)}}\color{orangered}{\left(\frac{\left(\frac{1+x}{1}\right)}{\left(\frac{1+x}{1}\right)}\right)}\]You can multiplying it like this, if that helps to see how the fractions will multiply out. That is the same as multiplying top and bottom by (1+x)

OpenStudy (anonymous):

@zepdrix what would the domain be?